OpenStudy (anonymous):

Can someone please help me? dividing powers 4. x^2y^9 divided by 2x^3y^6 5. a^3b^2c divided by a^3bc^5 6. 3pq^3 divided by 4p^4q^8r

OpenStudy (zzr0ck3r):

(x^2y^9)/(2x^3y^6) = x^(2-3)*y^(9-6)/2

OpenStudy (zzr0ck3r):

= y^3/(2x)

OpenStudy (anonymous):

can you do 5 & 6 please and thank you?

OpenStudy (zzr0ck3r):

all you need on all these is the fact that a^b/a^c = a^(b-c) = 1/a^(c-b)

OpenStudy (zzr0ck3r):

we are not here to do them for you, I showed you what I did, you attempt it and show me what your doing

OpenStudy (anonymous):

i know but i only have a few days to understand this and to get all of them done /: im stuck

OpenStudy (anonymous):

i'm still stuck on 5 & 6

OpenStudy (rajee_sam):

In all these types of problems you have to combine exponents with same base